# Causally Testing Gender Bias in LLMs: A Case Study on Occupational Bias

### v3

@@ Appendix A Gender Verbalizations
As can be seen in Table 2, task prompt number one uses a variety of expressions for different genders. Below is a complete list of expressions. Note that for all expressions, both probabilities of capitalized and non-capitalized expressions were measured and taken into account when computing probabilities of gender associations.
@@ Appendix B Occupation Data
We use occupation data from 2021 provided by the U.S. Bureau of Labor Statistics to obtain lists of jobs that are dominated by males and females. We did not use the twenty jobs with the highest ratio of male and females working in them each, as the data did contain highly specific job names that could better be summarized under umbrella terms. We therefore curated and summarized the data as well as possible. The resulting list of jobs with their corresponding ratios of males and females working in them can be found in Table 4.

### v4

@@ Appendix A Gender Verbalizations
As it can be seen in Table 2, task prompt number one uses a variety of expressions for different genders. Below is a complete list of expressions. Note that for all expressions, both probabilities of capitalized and non-capitalized expressions were measured and taken into account when computing probabilities of gender associations.
@@ Appendix B Occupation Data
We use occupation data from 2021 provided by the U.S. Bureau of Labor Statistics to obtain lists of jobs that are dominated by males and females. We did not use the twenty jobs with the highest ratio of males and females working in them each, as the data did contain highly specific job names that could better be summarized under umbrella terms. We therefore curated and summarized the data as well as possible. The resulting list of jobs with their corresponding ratios of males and females working in them can be found in Table 5.